### v2

# Measurement as Sheafification: Context, Logic, and Truth after Quantum Mechanics

Abstract
Quantum measurement is commonly posed as a dynamical tension between linear Schrödinger evolution and an ad hoc collapse rule. I argue that the deeper conflict is logical: quantum theory is inherently contextual, whereas the classical tradition presupposes a single global, Boolean valuation. Building on Bohr’s complementarity, the Einstein–Podolsky–Rosen argument and Bell’s theorem, I recast locality and completeness as the existence of a global section of a presheaf of value assignments over the category of measurement contexts. The absence of global sections expresses the impossibility of context-independent description, and Čech cohomology measures the resulting obstruction. The internal logic of the presheaf topos is intuitionistic, and the Ghose–Patra seven-valued contextual logic is exhibited as a finite Heyting algebra capturing patterns of truth, falsity and indeterminacy across incompatible contexts. Classical physics corresponds to the sheaf case, where compatible local data glue and Boolean logic is effectively restored. Measurement is therefore reinterpreted as sheafification of presheaf-valued truth rather than as a physical breakdown of unitarity. Finally, a $\sigma$$\lambda$ dynamics motivated by stochastic mechanics provides a continuous interpolation between strongly contextual and approximately classical regimes, dissolving the usual measurement paradoxes and apparent nonlocality as artefacts of an illegitimate demand for global truth.
Keywords: quantum measurement; contextuality; presheaves and sheaves; topos/Heyting logic; Saptabhaṅgī; $\sigma$$\lambda$ dynamics
1 The Classical Pact Between Logic and Physics
Classical mechanics admits a description in which physical quantities possess definite values simultaneously (like position and momentum) and independently of how they are measured. This feature aligns seamlessly with Boolean logic, where propositions are either true or false absolutely. Over time, this harmony hardened into a tacit pact: logic came to be regarded as universally valid and ontologically neutral, while physics merely instantiated it. The success of classical physics concealed the contingency of this arrangement. Logic appeared to precede and constrain physics, rather than to arise from the structure of physical description.
This harmony between classical physics and classical logic did not arise by historical accident alone. It received its most influential philosophical formulation in Kant’s response to Newtonian science [9]. Kant’s central insight was that the universal validity and necessity exhibited by Newtonian mechanics could not be derived from experience alone. Instead, such necessity must arise from the structures through which experience itself becomes possible.
According to Kant, space and time are not properties of things in themselves but forms of sensible intuition, while fundamental concepts such as object, causality, and substance are conditions under which phenomena can be thought at all. The laws of Newtonian physics thus appear universal and exceptionless because they describe a world already organised by these a priori forms. Crucially for our purposes, this organisation presupposes that physical systems possess determinate properties at each moment in time and that propositions about them admit unambiguous truth values.
In this way, Newtonian physics and classical logic were brought together within a single framework. Boolean logic, with its sharp distinction between true and false, was not merely compatible with classical mechanics; it appeared to be forced upon us by the very conditions of possible experience. The logical form of physical description was thus elevated from a contingent feature of successful theories to a seemingly universal necessity.
The enduring influence of this Kantian synthesis lies in its stability: once the logical preconditions of Newtonian science were identified with the preconditions of experience as such, the idea that alternative physical theories might call for alternative logical frameworks became effectively unthinkable. It is this deeply ingrained assumption—rather than any specific physical postulate— that quantum mechanics ultimately calls into question.
It is worth noting that even within classical physics this Kantian synthesis was not left entirely unchallenged. Einstein’s theory of relativity already undermined one of Kant’s central assumptions: the a priori status of Euclidean geometry. As Einstein himself emphasized in his Autobiographical Notes [10], the geometric structure of space is not fixed by the conditions of thought alone, but is subject to empirical determination by physical theory.
Yet despite this profound revision, relativity theory remained classical in a crucial sense. Physical states were still described globally by fields defined on spacetime, their evolution governed by deterministic equations, and physical propositions retained unambiguous truth values. What changed was not the logical form of physical description, but its geometric backdrop. Boolean logic survived relativity intact.
In hindsight, this illustrates an important distinction. Relativity revealed that geometry need not be a priori, but it left untouched the deeper assumption that physical reality admits a single, context-independent description. It is this remaining assumption—preserved by Einstein—that quantum mechanics would ultimately force us to abandon, much to Einstein’s discomfort.
2 Quantum Mechanics as a Crisis of Global Truth
Quantum mechanics marks a decisive break with the classical assumption that physical properties admit a single, context-independent description. This departure is already implicit in Bohr’s formulation of the Complementarity Principle [11], long before later no-go theorems made the point mathematically explicit. Bohr emphasized that the description of quantum phenomena depends ineliminably on the experimental conditions under which they are observed. The formalism itself assigns no meaning to properties in abstraction from such conditions.
Bohr’s use of the double-slit experiment to illustrate complementarity is especially telling. Under one experimental arrangement, designed to register interference, the quantum object must be described as a wave propagating through both slits. Under a mutually exclusive arrangement, designed to determine which slit the object passes through, the description is necessarily particle-like. These two descriptions are not merely practically incompatible; the conditions that make one meaningful physically preclude the other. There is no single experimental context in which both descriptions can be jointly applied. What is crucial is that each description is perfectly well-defined and objective within its own context, yet no global synthesis of the two is available.
Although Bohr did not formulate his insights in the language of formal logic, the implication is clear: the truth of propositions about quantum systems is conditioned by the experimental context. The attempt to ascribe simultaneous truth to all such propositions reproduces a classical demand that the theory itself does not support. Complementarity thus introduces contextuality into quantum mechanics at a foundational level, even without appealing to hidden variables or metaphysical assumptions.
This contextuality was later sharpened and formalized through results such as the Kochen–Specker theorem [12]. Unlike Bell-type arguments [13], which rule out local hidden-variable theories under assumptions of separability or locality, the Kochen–Specker theorem demonstrates the impossibility of assigning definite values to quantum observables in a way that is both non-contextual and consistent with the functional relations between observables. No assignment of pre-existing values can reproduce the predictions of quantum mechanics if those values are required to be independent of measurement context.
The notion of contextuality that emerges here differs in form from Bohr’s original discussion, but not in substance. In both cases, the obstruction lies in the attempt to maintain a globally valid assignment of truth values across mutually incompatible experimental arrangements. What Kochen–Specker makes explicit is that this obstruction is not a consequence of practical limitations or incomplete knowledge, but a structural feature of the theory itself.
Quantum mechanics thus confronts us with a new situation: propositions that are perfectly meaningful and decidable within a given context cannot, in general, be assembled into a single context-independent account. This failure of global truth is not pathological; it is intrinsic to the quantum description of nature. The persistence of the measurement problem and related paradoxes is a reflection of the continued use of classical, Boolean logic in a domain where its basic presuppositions no longer apply.
It is worth emphasising that several of the founders and early critics of quantum mechanics came close to recognising this impasse, yet stopped just short of its logical resolution. Bohr’s doctrine of complementarity correctly identified the indispensability of mutually incompatible experimental contexts, but retained a classical conception of logical description within each context without formalising the logical relations between them. The Kochen–Specker theorem went further in showing that no global assignment of definite values is possible, but it did so as a negative result: it demonstrated the impossibility of a classical valuation without replacing it by a new, explicitly contextual logic.
Von Neumann and Birkhoff made the boldest attempt to reform logic itself by introducing quantum logic, replacing Boolean lattices by the lattice of closed subspaces of Hilbert space [14]. However, this approach preserved a global, context-independent logical structure at a deeper level, and thereby shifted rather than resolved the underlying tension. Logical operations were altered, but the assumption of a single overarching logical space remained intact.
Reichenbach’s proposal of a three-valued probability logic likewise acknowledged the inadequacy of classical truth values, yet interpreted indeterminacy probabilistically rather than contextually [15]. As a result, it treated indefiniteness as a matter of partial truth or ignorance, rather than as a structural feature arising from incompatible descriptions.
What these influential approaches share is a reluctance to let logic itself be conditioned by physical context. Quantum mechanics, however, persistently forces such a conditioning. The ambiguity that remains in the measurement problem is not accidental; it is the residue of an unmet demand to rethink the very notion of truth in a context-dependent way. This essay proceeds from the claim that only by doing so can the conceptual tension at the foundation of quantum mechanics finally be dissolved.
The EPR Argument and the Demand for Global Truth
The tension between quantum mechanics and classical logic was brought into sharp focus by the celebrated Einstein–Podolsky–Rosen (EPR) argument [16]. EPR considered two systems that have interacted and then separated, so that their joint quantum state displays perfect correlations between suitably chosen observables. If one measures an observable $A$ on the first system, one can predict with certainty the outcome of a corresponding observable $B$ on the distant system, without in any way disturbing it. EPR then introduced a criterion of reality: if, without disturbing a system, one can predict with certainty the value of a physical quantity, then there exists an element of physical reality corresponding to that quantity.
Assuming locality in the strict sense that operations on the first system cannot instantaneously affect the second, EPR concluded that the second system must possess definite values for both of two incompatible observables, depending on which measurement is performed on the first. Since the quantum formalism assigns no such joint values to incompatible observables, EPR inferred that the quantum-mechanical description of physical reality is incomplete. In effect, their argument demands that, under the joint assumptions of locality and perfect correlation, there must exist a single, global assignment of values to all relevant observables: a global truth function extending across all measurement contexts.
Bohr’s immediate reply to EPR in 1935 made this contextual dependence completely explicit [17]. He grants that the two particles form a single entangled whole and that a measurement on one side allows one to predict with certainty the outcome of a suitably chosen measurement on the distant partner. But he denies that this justifies ascribing simultaneous reality to noncommuting quantities of the distant particle. The crucial point, for Bohr, is that the very meaning of “position of the second particle” or “momentum of the second particle” is fixed only within a well-defined experimental arrangement. The EPR argument illicitly combines facts obtained under mutually exclusive arrangements into a single global description. In Bohr’s language, there is indeed “no mechanical disturbance” of the distant system; rather, the choice of measurement here changes the conditions which define the possible predictions there. In modern terms, Bohr is already arguing that truth for quantum propositions is irreducibly contextual: one cannot form a single Boolean algebra of properties spanning all complementary setups at once, and the EPR demand for such a global truth assignment is therefore conceptually misplaced rather than empirically refuted. In contemporary language, this is precisely the demand that there should exist a global section of the presheaf of value assignments. EPR took the existence of such a global section to be a requirement of locality and completeness; the subsequent development of Bell’s theorem showed that no local hidden-variable theory can reproduce the full range of quantum correlations [13]. The usual conclusion is that one must either give up locality or accept some form of nonlocal “spooky action-at-a-distance.”
The perspective adopted in this paper is different. The EPR argument reveals not a failure of quantum mechanics to supply missing variables, but a clash between the presheaf-like, contextual structure enforced by the theory and the assumption that there must nonetheless exist a single global valuation compatible with locality. Once one acknowledges that quantum truth is intrinsically context-dependent and that the relevant semantic object is a presheaf without global section, the EPR demand for a local global truth assignment is seen to be ill-posed. The “nonlocality” that appears in the standard reading of Bell’s theorem is thus a symptom of insisting on a global logical structure that the physical theory does not support, rather than direct evidence for superluminal influences.
This way of phrasing the EPR and Bell arguments is closely related to the sheaf-theoretic analysis of non-locality and contextuality developed by Abramsky and Brandenburger [18], in which contextuality is likewise characterised as the obstruction to the existence of global sections of a presheaf of outcome assignments.
3 Why the Measurement Problem Is Conceptually Misplaced
Once the impossibility of global truth assignments has been recognised, the traditional framing of the measurement problem comes into focus as a misdiagnosis. The problem is conventionally posed as a demand for a physical mechanism that transforms quantum superpositions into definite classical outcomes. This demand, however, already presupposes that physical reality must at all times admit a single, global, context-independent description. It is this presupposition—rather than any deficiency in the quantum formalism—that gives rise to the appearance of paradox.
Classical logic embodies the assumption that propositions possess definite truth values independently of the conditions under which they are evaluated. In classical physics, this assumption is unproblematic: the theory admits a global phase space in which all observables have simultaneous, well-defined values. Logical operations such as conjunction and disjunction merely mirror this underlying structure. When classical logic is carried over uncritically into quantum theory, however, it enforces a requirement that the theory itself systematically refuses to satisfy. The insistence on global truth becomes an external constraint imposed on a fundamentally contextual framework.
The resulting conflict inevitably manifests as a physical catastrophe. If the quantum state is taken to evolve unitarily at all times, classical logic demands that measurement outcomes nevertheless be globally definite. This forces the introduction of an additional dynamical process—collapse—whose sole function is to restore a logical condition that the unitary dynamics violates. Collapse thus appears not as a physical necessity arising from the theory, but as a logical repair mechanism introduced to save a classical notion of truth.
This tension was already sharply perceived by Schrödinger in his 1935 analysis of the ‘present situation in quantum mechanics’ [19]. Starting from the linear, deterministic Schrödinger evolution, he stresses that, if the formalism is taken at face value, microscopic superpositions are unavoidably amplified into grotesque macroscopic superpositions—live cat plus dead cat. At the same time, he notes that the standard account prescribes a suspension of this unitary evolution during measurement, replacing it by an abrupt, probabilistic “jump” to a definite outcome, without offering any precise dynamical law for when and how this happens. Schrödinger thus treats the collapse postulate less as a physical mechanism than as a sign of conceptual deficiency: a rule introduced purely to reconcile the formalism with the definite character of experience. In the sheaf-theoretic reading proposed here, this tension is not resolved by modifying the Schrödinger dynamics, but by recognising that the passage from entangled, context-dependent descriptions to classically communicable outcomes is a change in the logical regime—a sheafification of presheaf-valued truth—rather than a mysterious violation of the unitary law.
Seen in this light, the familiar paradoxes of quantum mechanics acquire a new interpretation. Schrödinger’s cat, Wigner’s friend, and related scenarios do not signal a breakdown of physical law; they reveal the incompatibility between a contextual theory and a non-contextual logic. Each paradox arises when one attempts to combine descriptions that are valid within different contexts into a single global account. The contradiction is logical before it is physical.
Interpretational strategies respond to this tension in characteristic ways. Some, such as objective collapse models, modify the dynamics to enforce global definiteness. Others, such as many-worlds approaches, preserve unitary evolution at the cost of multiplying classical realities. Still others introduce privileged observers or appeal to consciousness. What these diverse responses have in common is their attempt to secure classical logical absoluteness rather than to question it. They all treat context-dependence as an anomaly to be eliminated, rather than as a structural feature to be accommodated.
The central claim of this essay is that this strategy is misguided. Quantum mechanics does not demand the explanation of a mysterious physical transition, but the abandonment of an inappropriate logical expectation. Once truth itself is recognised as context-dependent, the demand for a universal, collapse-inducing mechanism loses its force. The measurement problem, properly understood, is not a problem that calls for a new physical process, but a symptom of insisting on a form of logic that quantum theory no longer supports.
It is worth noting in this connection that Einstein complained in a letter written on 19 June, 1935 to Schrödinger that the EPR paper, drafted by Podolsky “for reasons of language” had allowed the main point to be “smothered by formalism (Gelehrsamkeit)” [20]. In his later reflections, he distanced his concern from specific hidden-variable programmes and suggested instead that the difficulty lay deeper than the addition of further physical parameters. It is therefore clear that Einstein felt profoundly that something essential was missing in quantum mechanics, even though he did not—and perhaps could not have— anticipated that the missing element would lie in the logical structure through which the theory is interpreted.
4 Context as the Primitive: From Measurement Setups to Categories
If the measurement problem is really a problem of logic rather than of dynamics, we must ask what the basic logical units are. In the present proposal the primitive notion is that of a measurement context. Roughly speaking, a context is a physically real experimental arrangement: a choice of observables, an arrangement of apparatus, a temporal ordering of interventions, and an environment in which the standard quantum predictions can be applied without further qualification.
Two features of such contexts are crucial. First, within any given context, the usual quantum formalism delivers definite probabilistic predictions; once an outcome is registered, it can be reported as an unambiguous fact. Second, different contexts may be mutually incompatible: the conditions that make one arrangement meaningful may exclude another. Bohr already insisted on this point in his discussions of complementary experiments; the double-slit arrangement designed to reveal interference excludes, as a matter of physical principle, the arrangement that would reveal which-path information. There is no meta-context in which both descriptions apply simultaneously.
Mathematically, it is natural to organise such contexts into a category $C$. The objects of $C$ are measurement contexts. A morphism $f:C\to D$ is interpreted as a physically meaningful refinement or coarse-graining: passing from $C$ to $D$ by adding compatible observables, increasing resolution, or otherwise sharpening the description. Composition corresponds to performing such refinements in succession, and identity arrows represent leaving a context unchanged. Importantly, not every pair of contexts need be related by a morphism: the absence of an arrow $C_{1}\to C_{2}$ records physical incompatibility rather than logical contradiction.
Within each context $C$ one may consider the propositions about the system that are testable in $C$ and the corresponding truth values assigned by the theory. Classically this data would be packaged into a single global phase space; in the present viewpoint it is deliberately kept local to each context. The question of how such local assignments behave when one passes from one context to another, and whether they can be glued into a single global description, is then expressed in the language of functors on $C$.
This is precisely where presheaves enter. A presheaf on $C$ assigns to each context $C$ the appropriate set of truth assignments, expectation values, or outcome structures, together with restriction maps along morphisms in $C$. The next section will show how presheaves provide the natural semantic framework for context-dependent quantum truth, and how the failure of a global section expresses, in a mathematically precise way, the breakdown of classical, context-independent description.
5 Presheaves as the Natural Semantics of Quantum Truth
The conceptual shift required by quantum mechanics is most clearly expressed once attention is moved away from absolute states and toward relations between contexts. Category theory [21] (see also Appendix A for a concise mathematical introduction) provides a language precisely suited for this purpose, as it allows one to describe not only objects, but also the web of relations between them. Related categorical reconstructions of quantum mechanics have been developed by Abramsky and Coecke [22, 23], who axiomatise the theory at the level of strongly compact closed (or dagger compact) categories in order to capture the compositional structure of quantum processes and information flow. Their “categorical quantum mechanics” programme shows that a large part of the Hilbert-space formalism can be recovered from purely structural assumptions on such process categories. The use of category theory in the present work is complementary to this: rather than reaxiomatising the state–process calculus itself, it organises measurement contexts into a category and uses presheaves and cohomological tools to analyse the logical and contextual structure of truth and measurement. In the present setting, the objects of interest are measurement contexts, understood as physically real experimental arrangements, while the relations describe how one context may refine or extend another. Importantly, not all contexts need be mutually related: incompatibility is expressed not by contradiction, but by the absence of a relation.
A category, in this intuitive sense, is nothing more than a structured collection of contexts together with the physically meaningful ways of passing from one to another. What matters is not the internal composition of each context in isolation, but how descriptions change when the context is varied. This relational viewpoint already departs from classical thinking, where a single global domain is presupposed from the outset.
Presheaves arise naturally once one asks how truth or physical description is assigned relative to contexts. A presheaf associates to each context the set of statements, values, or assignments that are meaningful within that context, and specifies how these assignments are restricted when passing to a more limited or coarser context. Crucially, presheaves make no demand that locally valid descriptions fit together into a single global picture. They formalise the possibility that truth may be well-defined within each context, yet resist unification across incompatible ones.
This feature makes presheaves particularly well suited to quantum mechanics. The theory consistently provides context-dependent descriptions—expectation values, probabilities, or definite outcomes—while denying the existence of a single context in which all such descriptions can be simultaneously realised. The failure of global truth in quantum mechanics is therefore not an anomaly, but exactly the behaviour presheaves are designed to accommodate.
Sheaves represent a special case within this framework. A sheaf imposes a stronger consistency requirement: whenever local descriptions agree on all overlaps between contexts, they must arise from a unique global description. This is precisely the logical structure tacitly assumed in classical physics. From this perspective, classical logic emerges not as a fundamental principle, but as a consequence of working in regimes where the sheaf condition happens to be satisfied.
The distinction between presheaves and sheaves thus mirrors the distinction between quantum and classical truth. Quantum theory naturally gives rise to presheaf-like semantics, while classical physics corresponds to the special case in which presheaf data can be glued into global sections.
Cohomology provides a further conceptual refinement. Rather than merely stating that global truth may fail, cohomological tools offer a way to characterise and measure this failure. When locally valid assignments cannot be consistently glued together, the obstruction can be represented by cohomology classes. In physical terms, nontrivial cohomology signals the presence of irreducible contextuality, while the vanishing of such obstructions corresponds to the emergence of classical, globally consistent descriptions.
In this way, categories encode contexts, presheaves encode context-dependent truth, sheaves encode classical consistency, and cohomology records the precise sense in which quantum descriptions resist global unification. These ideas will provide the conceptual foundation for the more detailed analysis that follows.
Intuitionistic Logic Inside Presheaves
The logical behaviour of presheaves already departs from the classical, two-valued picture in a fundamental way. The internal logic of any category of presheaves is not Boolean but intuitionistic. In classical logic, propositions are assumed to be either true or false absolutely, and the principles of excluded middle ($P\vee\neg P$) and double negation ($\neg\neg P\Rightarrow P$) hold universally. Intuitionistic logic abandons these principles as general laws. A statement is taken to be true only when a suitable construction or justification is available, and it need not be the case that either $P$ or $\neg P$ holds in the absence of such a construction.
Presheaves give a natural semantic home for intuitionistic logic because truth is evaluated relative to context. A proposition about a quantum system may have a well-defined truth value in one measurement context, a different value in another, and no determinate value at all in incompatible contexts. The collection of all such truth values in a presheaf topos forms a Heyting algebra rather than a Boolean algebra: it supports conjunction, disjunction and implication, but not in a way that forces every proposition to be either globally true or globally false. Intuitively, truth can grow monotonically as contexts are refined, and the logical structure records this possibility.
Seen in this light, classical Boolean logic appears as a special, degenerate case of intuitionistic logic, recovered precisely when the presheaf data collapse to a single global section. In that regime the Heyting algebra of truth values becomes Boolean: excluded middle and double negation are restored, and propositions behave as if they had context-independent truth values. Outside this regime, however, the more flexible, context-sensitive structure of intuitionistic logic is indispensable.
Contextual Multi-Valued Logic
This intuitionistic background naturally paves the way for a multi-valued treatment of quantum propositions. Once truth is recognised as context-dependent, it is no longer adequate to speak only of “true” and “false” simpliciter. One must also distinguish cases in which a proposition is true in one context and false in another, or true in one context and indeterminate in another, and so on. These patterns of contextual variation themselves become the relevant “truth values.”
The seven-valued scheme proposed by Ghose and Patra [24] can be viewed in precisely this way (see Appendix B for some technical details). Each of the seven values—“true,” “false,” “indeterminate,” and the four mixed cases—represents a distinct mode in which truth may vary across incompatible contexts. Formally, these seven values can be organised into a finite Heyting algebra: they admit logical operations of conjunction, disjunction and implication, but do not collapse to a simple two-valued Boolean structure. In this sense, the GP logic may be regarded as a concrete, context-sensitive instance of intuitionistic semantics tailored to quantum phenomena.
What distinguishes the GP scheme from more abstract versions of intuitionistic logic is its explicitly contextual and relational character. The multi-valuedness does not arise from a vague notion of partial truth or degree of belief, but from the way propositions about quantum systems take on different truth values in physically incompatible measurement arrangements. The resulting structure is therefore not a mere generalisation of classical logic, but a logical reflection of the contextual architecture that presheaves and their intuitionistic logic make precise.
6 Sheaves and the Emergence of Classical Logic
A sheaf is a presheaf with the added property that compatible local data uniquely determine a global section. Sheaves thus encode the logical structure presupposed by classical physics: context-independent truth and Boolean reasoning. Classical logic is not fundamental in this view; it emerges when physical circumstances permit gluing. Classicality is therefore a special regime, characterised by the satisfaction of sheaf conditions.
In the next subsection, this emergent view will be applied directly to the measurement process.
Measurement Reinterpreted: Sheafification, Not Physical Collapse
From this perspective, the traditional problem of measurement acquires a very different character. In the interpretation associated with Bohr, there is no measurement problem in the later, technical sense. Bohr consistently rejected the demand for a physical account of wavefunction collapse and instead emphasised the necessity of describing experimental outcomes in a classical language refined through the development of Newtonian mechanics and Maxwellian electrodynamics. What mattered for him was not the ontological status of the quantum state, but the possibility of unambiguous communication of experimental results.
This insistence on classical description may be understood, in modern terms, as a demand for global consistency. Classical language presupposes that outcomes can be reported as definite, shared facts, independent of the particular experimental context in which they are obtained. In the present framework, this corresponds precisely to the existence of global sections—descriptions that are valid across all relevant contexts and can therefore be communicated without ambiguity. Bohr’s requirement thus amounts to privileging those descriptions that satisfy a sheaf-like condition.
What Bohr carefully avoided was the attribution of such global descriptions to quantum systems prior to measurement. Quantum theory, in his view, provides only context-dependent accounts tied to specific experimental arrangements. Measurement does not reveal a pre-existing global state; it marks the point at which a description must be rendered in classical terms in order to be communicable. In this sense, Bohr’s “cut” separates presheaf-level quantum descriptions from their sheaf-level classical articulation.
Reinterpreted in this way, measurement need not be understood as a physical collapse interrupting unitary dynamics. Rather, it is a logical transition from contextual, presheaf-based descriptions to globally consistent, sheaf-like ones. The apparent discontinuity arises not in the underlying dynamics, but in the imposition of a requirement of global truth. The collapse postulate enters only when this logical transition is misread as a physical process.
Seen through the lens of sheaf theory, Bohr’s insistence on classical language appears not as an ad hoc philosophical restriction, but as an implicit recognition that communication itself requires globalisation. The measurement problem emerges only when this requirement is demanded of quantum descriptions themselves, rather than of the language in which their outcomes are reported.
7 Dynamics as a Continuous Passage Between Logical Regimes
Up to this point the discussion has been largely structural: quantum theory was seen to enforce a presheaf-like, context-dependent notion of truth, while classical physics corresponded to the special case in which those presheaf data can be glued into a sheaf of globally valid descriptions. To make this picture physically credible, one needs a dynamical mechanism that allows for a continuous passage between these two regimes. Nelson’s stochastic mechanics provides a natural starting point [25].
Nelson begins from the classical Newtonian picture, but assumes that particles undergo a Brownian motion with a diffusion constant $\sigma$ superposed on their regular motion. Instead of a single velocity field, one has forward and backward mean velocities, and the particle trajectories are described by stochastic differential equations. The key result is that, under suitable assumptions, the ensemble dynamics of such stochastic trajectories can be recast into the familiar Schrödinger equation. The Planck constant $\hbar$ then appears not as a primitive quantum postulate, but as a parameter fixed by the relation
$\hbar=m\sigma,$
where $m$ is the mass and $\sigma$ is the diffusion parameter characterising the underlying Brownian motion. [^1] More standardly one writes $\nu=\hbar/(2m)$ for the diffusion constant; the present notation absorbs numerical factors into $\sigma$ for simplicity.
This identification has an important conceptual consequence. It shows that what is usually regarded as a fundamental constant can be viewed, at the level of the underlying stochastic processes, as a compound parameter linking mass and diffusion. In particular, one may consider variations of $m$ and $\sigma$ that leave $\hbar$ fixed, or, more generally, explore regimes in which the effective strength of the underlying diffusion is altered. The degree to which the resulting dynamics exhibits characteristic quantum features can then be controlled continuously.
A complementary way to express this continuous control is to work with the hydrodynamic form of the Schrödinger equation. Writing the wave function in polar form,
$\psi=\sqrt{\rho}\,e^{iS/\hbar},$
and separating real and imaginary parts, one obtains a continuity equation for the probability density $\rho$ and a modified Hamilton–Jacobi equation for the phase $S$:
$\frac{\partial S}{\partial t}+\frac{(\nabla S)^{2}}{2m}+V+Q=0,$
where $V$ is the classical potential and
$Q=-\,\frac{\hbar^{2}}{2m}\,\frac{\nabla^{2}\sqrt{\rho}}{\sqrt{\rho}}$
is the so-called quantum potential. Setting $Q=0$ recovers the classical Hamilton–Jacobi equation; retaining $Q$ yields the full quantum dynamics.
Rosen [26] observed that the classical Hamilton–Jacobi and continuity equations can themselves be combined into a complex “classical Schrödinger equation” in which the quantum potential term is absent. From this viewpoint, the essential difference between classical and quantum dynamics is precisely the presence or absence of $Q$. This suggests introducing a one-parameter family of interpolating dynamics by writing
$\frac{\partial S}{\partial t}+\frac{(\nabla S)^{2}}{2m}+V+\lambda\,Q=0,$
with $0\leq\lambda\leq 1$ [27]. The case $\lambda=1$ reproduces the usual quantum dynamics, while $\lambda=0$ yields the classical Hamilton–Jacobi equation. Intermediate values of $\lambda$ describe regimes in which the influence of the quantum potential is progressively suppressed, and the dynamics moves continuously from quantum-like to classical-like behaviour.
In the stochastic picture, the parameter $\lambda$ can be regarded as a function of the underlying diffusion, $\lambda=\lambda(\sigma)$, encoding how strongly the quantum potential survives coarse-graining and environmental influence. Large diffusion (or suitable choices of $m$ and $\sigma$ satisfying $\hbar=m\sigma$) support $\lambda\approx 1$ and hence fully quantum behaviour; as the effective diffusion weakens, the quantum potential term is diminished and $\lambda$ tends toward 0. The $\sigma$$\lambda$ dynamics thus implements a continuous deformation of the underlying Hamilton–Jacobi structure between quantum and classical limits.
Crucially, this continuous dynamical interpolation can be read as a continuous passage between logical regimes. In the $\lambda\approx 1$ regime, where the quantum potential is fully active, trajectories are highly sensitive to contextual information, and the space of measurement contexts exhibits strong nontriviality. This is the domain in which presheaf semantics and intuitionistic, multi-valued logic are indispensable, and cohomological obstructions to global truth are generically present. As $\lambda$ decreases and the quantum potential is progressively suppressed, trajectories approach classical behaviour, incompatibilities between contexts lose their operational significance, and the presheaf of truth values becomes increasingly close to a sheaf. In the limit $\lambda\to 0$, the cohomological obstructions vanish, a global section emerges, and classical Boolean logic is effectively restored.
In this way, the $\sigma$$\lambda$ dynamics does more than interpolate between two mathematical equations, or even between two physical regimes. It realises a continuous transition from a world in which truth is irreducibly contextual and presheaf-based, to one in which truth can be treated, to excellent approximation, as global and classical. Measurement, understood as the selection of a sheaf-like description, is thereby anchored in a dynamical process rather than in a postulated discontinuity.
8 Logic as Emergent and the Dissolution of the Measurement Problem
We are now in a position to look back and reconsider what has been at stake in the measurement problem. The traditional formulation takes for granted that there is a single, context-independent logical framework in which all physical propositions must be evaluated. Within that framework, the coexistence of unitary evolution and collapse appears paradoxical. Unitary dynamics spreads possibilities into superpositions; collapse restores definiteness in a way that seems to violate the very principles governing the evolution between measurements. The problem, as usually stated, is how to reconcile these two modes of description within one physical theory.
The route taken in this essay is different from the standard treatments. Rather than starting from the quantum formalism and asking how it might be made to fit within a fixed classical logical framework, we have asked what kind of logical framework the formalism itself suggests. In the perspective developed here, quantum mechanics does not undermine the idea of a coherent physical world, but it does challenge the assumption that this world must always admit a single, context-independent global description. The category of measurement contexts, and the presheaves defined over it, provide a mathematically precise language for articulating this re-interpretation. Quantum theory furnishes context-dependent descriptions that are internally consistent within each context, yet resist amalgamation into a single global picture. Classical physics, by contrast, corresponds to the special case in which these presheaf data satisfy a sheaf condition, admitting global sections and thereby supporting a Boolean notion of truth.
From this perspective, logic is no longer a fixed background against which physics unfolds, but an emergent structure reflecting the way physical contexts are organised. In strongly quantum regimes, where contextuality is ineliminable and cohomological obstructions to global truth are present, the appropriate internal logic is intuitionistic and multi-valued, exemplified by context-sensitive schemes such as the GP seven-valued logic. In classical regimes, where the obstructions vanish and global sections exist, this richer logical structure collapses to the familiar two-valued Boolean form. Logic, in other words, is conditioned by the physical organisation of contexts: it is presheaf-like and contextual when the world forces contextuality upon us, and effectively Boolean when the world allows globalisation.
The $\sigma$$\lambda$ dynamics gives this picture a continuous physical realisation. By interpreting the quantum potential as a $\lambda$-scaled term interpolating between quantum and classical Hamilton–Jacobi dynamics, and linking $\lambda$ to the underlying stochastic parameter $\sigma$, one obtains a smooth passage from a regime in which contextual effects are dominant to one in which they are negligible. As $\lambda$ moves from $1$ toward 0, the influence of the quantum potential fades, the operational distinction between incompatible contexts diminishes, and the presheaf of truth values becomes increasingly close to a sheaf. In the limit, cohomological obstructions disappear and a global, approximately classical description emerges. The transition from quantum to classical is thus not a sudden jump, but a deformation of both dynamics and logic.
In this light, the measurement problem loses its air of mystery. What had appeared as a physical catastrophe—the abrupt collapse of the wavefunction—can now be understood as a change in the admissible form of description. Measurement corresponds to the selection, in practice, of sheaf-like, globally communicable accounts from an underlying presheaf of contextual quantum possibilities. Collapse is not a fundamental physical process layered on top of unitary evolution, but a logical projection associated with the demand for classical, context-independent reports of experimental outcomes. The paradox arises only if one insists that the presheaf-level quantum description must itself obey classical logical constraints.
Thus the measurement problem is not so much solved as dissolved. Once logic is recognised as emergent from the physical organisation of contexts, rather than imposed upon it, there is no longer a need to postulate an ad hoc collapse mechanism or to invoke observers or consciousness as external agents. The two “processes” of von Neumann are reinterpreted as two regimes of description: a presheaf-like, contextual regime appropriate to quantum systems, and a sheaf-like, Boolean regime appropriate to classical communication and macroscopic experience. The task is not to reconcile incompatible dynamics within a fixed logical frame, but to understand how different logical structures arise from the same underlying physical theory.
9 A Historical Afterword: Contextual Logic and Early Anticipations
The view developed in this essay has been motivated entirely by the internal demands of quantum theory: the breakdown of global truth, the centrality of measurement contexts, and the need for a logical framework in which context-dependence is not an anomaly but a structural feature. It is therefore striking, though historically accidental, that closely related ideas appear in a very different setting within classical Indian philosophy.
The Jaina doctrine of sevenfold predication, or Saptabhaṅgī, is often presented as a refinement of the broader principle of syādvāda (the “doctrine of may-be”) [28]. Instead of assigning a single, context-independent truth value to a proposition, the Jaina logicians distinguished seven possible predications: “in some respect, it is”; “in some respect, it is not”; “in some respect, it both is and is not”; “in some respect, it is indescribable”; and three further mixed combinations involving indescribability. The crucial point is that each predication is explicitly indexed by the qualifier syāt (“in some respect”): truth and falsity are not taken as absolute, but as relative to a standpoint, viewpoint, or mode of consideration.
In its original setting, this multi-valued, standpoint-dependent logic served primarily metaphysical and epistemological aims. It was intended to reconcile apparently incompatible perspectives on a complex reality without collapsing them into a single, privileged view. There is no suggestion of Hilbert spaces, measurement operators, or stochastic dynamics. Nevertheless, the formal pattern is recognisably that of a contextual logic: the same proposition may receive different evaluations under different, mutually irreducible perspectives, and the task of logic is to classify these evaluative patterns rather than to eliminate them.
The contextual seven-valued scheme developed by Ghose and Patra transposes this basic idea into a quantum setting. The seven values are no longer vague modalities of assertion, but precise patterns of truth, falsity, and indeterminacy across incompatible measurement contexts. Formally, as argued earlier, they can be represented as subobjects in a presheaf topos over the category of contexts, and their structure is that of a finite Heyting algebra rather than a Boolean one. In this sense, the GP logic may be regarded as a concrete, physically motivated instance of the kind of intuitionistic, context-sensitive semantics that presheaves naturally support.
It would be a mistake, however, to present the Jaina tradition as a “source” of quantum logic, or to suggest that quantum mechanics has been anticipated in any straightforward way by Saptabhaṅgī. The historical development of quantum theory is independent, and the mathematical tools employed here are entirely modern. The point of recalling the Jaina doctrine is more modest and, perhaps, more interesting: it shows that the idea of a reality that admits only context-dependent descriptions, and of a logic that classifies such descriptions without forcing them into a single global frame, is not an alien intrusion into human thought. Quantum theory compels us, for strictly physical reasons, to rediscover a possibility that had already been explored, in another guise, in a very different context and in a very different intellectual tradition.
10 Summary and Outlook
The analysis developed in this paper has suggested that the measurement problem in quantum mechanics is not primarily a dynamical paradox, but a manifestation of a deeper logical tension. The standard formulation tacitly presupposes that physical reality must be describable by a single, context-independent, globally valid Boolean valuation. Quantum theory, by contrast, forces upon us a world in which experimental arrangements are mutually incompatible and in which no global assignment of sharp values to all observables is possible. The apparent conflict between unitary evolution and wavefunction collapse, and the familiar worries about nonlocal “spooky action-at-a-distance”, are symptoms of this mismatch between a classical logical ideal and the contextual structure imposed by the quantum formalism itself.
By making measurement contexts explicit and organising them into a category, we have seen that presheaves provide a natural semantics for quantum truth. Context-dependent value assignments form presheaves on the category of contexts; the failure of a global section becomes a precise expression of the breakdown of classical, context-independent description. The internal logic of the resulting presheaf topos is intuitionistic rather than Boolean, and concrete schemes such as the Ghose–Patra seven-valued logic can be understood as finite Heyting algebras encoding distinct patterns of truth, falsity and indeterminacy across incompatible contexts. Sheaves then mark the regime in which local data can be glued into global descriptions: classical physics appears as the special case in which presheaf data satisfy a sheaf condition and Boolean logic is effectively restored.
On this background, the traditional postulate of collapse can be reinterpreted as a logical projection rather than a mysterious physical process. Bohr’s insistence on classical language is seen as a demand for sheaf-like, globally communicable descriptions, while the EPR argument and Bell’s theorem are recognised as attempts to impose a global truth assignment where only contextual presheaf data are available. Schrödinger’s unease about the coexistence of linear evolution and its suspension during measurement is thus read not as evidence for a physical discontinuity, but as an early recognition of a mismatch between the formalism and the classical logical frame into which it was being forced.
The $\sigma$$\lambda$ dynamics provides a physical realisation of the continuous passage between these logical regimes. By interpreting the quantum potential as a $\lambda$-scaled modification of the Hamilton–Jacobi dynamics, linked to an underlying stochastic parameter $\sigma$, one obtains a smooth interpolation from strongly contextual quantum behaviour to approximately classical behaviour in which cohomological obstructions vanish and global sections emerge. The dissolution of the measurement problem and the disappearance of “nonlocality” therefore coincide: both arise from abandoning the requirement of a single global Boolean description in favour of a contextual, presheaf-based semantics that is sheafified only where the physical organisation of contexts permits it.
Quantum gravity and the sheafification of spacetime.
A natural question is how the present perspective bears on quantum gravity. In the canonical approach, the central condition is the Wheeler–DeWitt constraint, $\widehat{\mathcal{H}}\Psi=0$, understood as the quantization of the Hamiltonian constraint of general relativity [29]. Although this is a manifestly quantum equation, its semiclassical (Born–Oppenheimer/WKB) expansion provides a standard bridge to classical spacetime: for a wavefunctional of the form $\Psi[h]\sim A[h]\exp(iS[h]/\hbar)$, the leading order yields the Einstein–Hamilton–Jacobi equation for $S[h]$, reproducing classical Einstein dynamics for a chosen geometric branch, with higher orders describing quantum fields on that background [30, 31]. From the present viewpoint, the recovery of classical general relativity may therefore be read as a two-step emergence: first, restriction to an appropriate quasi-classical branch within the constraint-satisfying sector; second, restoration of descent so that locally defined geometric data (chart, frame or connection descriptions on overlapping regions) glue into a globally communicable spacetime geometry. This use of gluing as a diagnostic of “classicality” resonates with earlier proposals to bring topos and sheaf ideas to quantum theory and quantum gravity [32] and with finitary sheaf approaches aimed at approximating continuum spacetime structure [33]. In our setting, the distinctive claim is that such restoration of global gluing can be tied to a continuous classicalization controlled by the $\sigma$$\lambda$ dynamics.
Several directions for further work suggest themselves. On the conceptual side, one may seek a more systematic classification of contextual logics within presheaf topoi and a clearer comparison with other approaches to quantum logic, including the Birkhoff–von Neumann programme and various topos-theoretic reconstructions. On the mathematical side, it would be natural to refine the cohomological analysis of contextuality and to explore more fully how the $\sigma$$\lambda$ dynamics controls the transition between nontrivial and trivial cohomology. On the physical side, one may study concrete models—for instance, simple interferometric or spin systems—in which the deformation from presheaf-like to sheaf-like behaviour can be analysed quantitatively, and investigate whether intermediate regimes admit experimental signatures. Whatever the outcome of such developments, the central lesson remains: logic is not a neutral backdrop for physics, but an emergent structure reflecting the way our physical world organises its contexts.
11 Appendix A: Introduction to Category Theory
Having given an intuitive account of the main thesis being presented here, a slightly more technical account follows for the interested reader. Category theory is often described as the mathematics of structure and relationships. Instead of focusing on elements inside sets or algebraic structures, it studies objects and arrows (morphisms) between them.
A category $\mathcal{C}$ consists of:
  • objects $A,B,C,\dots$
  • morphisms (arrows) $f:A\to B$
  • composition: if $f:A\to B$ and $g:B\to C$, then $g\circ f:A\to C$
  • identity arrows: for each object $A$, an arrow $\mathrm{id}_{A}:A\to A$
satisfying:
  • associativity: $(h\circ g)\circ f=h\circ(g\circ f)$
  • unit laws: $\mathrm{id}_{B}\circ f=f=f\circ\mathrm{id}_{A}$
Despite its simplicity, this structure is powerful enough to unify algebra, topology, logic, computation, and quantum theory.
Some basic examples are sets (objects are sets and morphisms are functions). groups (objects are groups and morphisms are group homomorphisms), topological spaces (objects are spaces and morphisms are continuous maps), Hilbert spaces (objects are Hilbert spaces and morphisms are bounded linear maps) and posets.
Category theory tells us:
  • what structures preserve what structures (via morphisms);
  • how structures compose;
  • when two structures should be considered the same (up to categorical equivalence rather than elementwise equality).
A functor $F:\mathcal{C}\to\mathcal{D}$ assigns:
  • to each object $A$ an object $F(A)$,
  • to each arrow $f:A\to B$ an arrow $F(f):F(A)\to F(B)$,
preserving identities and composition.
Examples include:
  • the forgetful functor $\textbf{Grp}\to\textbf{Set}$,
  • the free functor $\textbf{Set}\to\textbf{Grp}$,
  • homology functors in algebraic topology.
Natural Transformations
Given functors $F,G:\mathcal{C}\to\mathcal{D}$, a natural transformation $\eta:F\Rightarrow G$ assigns to each object $A$ a morphism $\eta_{A}:F(A)\to G(A)$ such that for every arrow $f:A\to B$,
$G(f)\circ\eta_{A}=\eta_{B}\circ F(f).$
Natural transformations express canonical, structure-preserving comparisons between functors.
A construction such as a product, coproduct, limit, colimit, kernel, etc. is defined not by its internal elements but by how it relates to other objects.
Universal properties ensure uniqueness up to unique isomorphism.
Adjunctions
Functors $F:\mathcal{C}\to\mathcal{D}$ and $G:\mathcal{D}\to\mathcal{C}$ form an adjoint pair, written $F\dashv G$, if there is a natural isomorphism:
$\mathrm{Hom}_{\mathcal{D}}(F(X),Y)\cong\mathrm{Hom}_{\mathcal{C}}(X,G(Y)).$
Adjunctions explain:
  • free/forgetful constructions,
  • Galois connections,
  • the categorical meaning of quantifiers,
  • dualities and fundamental constructions in algebra and topology.
Let $\mathcal{C}$ be a category of contexts. A presheaf on $\mathcal{C}$ is a functor
$F:\mathcal{C}^{op}\to\mathbf{Set}.$
It assigns:
  • a set $F(C)$ of data available in context $C$;
  • for every refinement $f:C\to D$, a restriction map
    $F(f):F(D)\to F(C).$
Physically:
  • Context $C$ may be a measurement setting, coarse-graining scale, reference frame, or $\sigma$$\lambda$ diffusion regime.
  • $F(C)$ is the set of observables, trajectories, or field-values accessible in context $C$.
  • $F(f)$ expresses how data in a finer context restricts to a coarser one.
Presheaves impose no global consistency. They naturally model contextuality and quantum-like behaviour.
Sheaves: Gluing of Local Data
A presheaf $F$ is a sheaf if compatible data on overlapping contexts can be uniquely glued into a global piece of data.
Given a cover $\{U_{i}\}$ of $U$:
$s_{i}\in F(U_{i}),\quad s_{i}|_{U_{i}\cap U_{j}}=s_{j}|_{U_{i}\cap U_{j}}$
implies
  • existence: a global $s\in F(U)$ with $s|_{U_{i}}=s_{i}$,
  • uniqueness: $s$ is unique.
Physically:
  • Classical fields behave as sheaves.
  • Classical probability distributions behave as sheaves.
  • Classical limits of $\sigma$$\lambda$ dynamics (as $\lambda\to 0$) approach sheaf-like behaviour.
Why Quantum Physics is Presheaf-Based
Quantum systems violate the gluing condition. Local sections exist, but global sections do not:
$\Gamma(F)=\varnothing.$
This is the mathematical signature of contextuality.
Sheafification = Classicalization
For any presheaf $F$, there is a canonical sheafification $F^{\sharp}$, which forces gluing by construction. Physically:
$\text{presheaf}\quad\text{(quantum-like)}\quad\longrightarrow\quad\text{sheaf}\quad\text{(classical)}.$
Cohomology: What It Measures
Given a presheaf $F$, the first Čech cohomology group
$\check{H}^{1}(\mathcal{C},F)$
measures the obstruction to gluing local sections into a global one.
A Čech-cohomological Obstruction to Contextuality
Let $\{C_{i}\to C\}_{i\in I}$ be a family of (overlapping) contexts covering a given context $C$ (e.g. the maximal compatible measurement contexts of an experimental scenario). Write $C_{ij}:=C_{i}\cap C_{j}$ and $C_{ijk}:=C_{i}\cap C_{j}\cap C_{k}$ for pairwise and triple overlaps.
To speak of Čech cohomology one works with an abelian coefficient presheaf. A convenient choice—used explicitly in cohomological contextuality—is to pass from a Set-valued presheaf $F$ to the presheaf of free abelian groups $\mathbb{Z}[F]$ generated by its sections, so that formal differences of restricted sections are meaningful [34].
A choice of local data is a Čech 0-cochain
$s=(s_{i})_{i\in I},\qquad s_{i}\in F(C_{i}).$
Its coboundary is the $1$-cochain $(\delta s)=(\delta s)_{ij}$ given on overlaps by
$(\delta s)_{ij}\;:=\;s_{j}|_{C_{ij}}\;-\;s_{i}|_{C_{ij}}\;\in\;\mathbb{Z}[F](C_{ij}).$
The sheaf gluing condition is precisely $\delta s=0$ (pairwise agreement on overlaps), in which case the family $\{s_{i}\}$ pastes to a global section $s\in F(C)$.
When $\delta s\neq 0$, the local sections fail to glue. Since $\delta^{1}\!\circ\delta^{0}=0$, the cochain $\delta s$ is automatically a $1$-cocycle, and its cohomology class
$\delta s]\;\in\;\check{H}^{1}(\{C_{i}\},\mathbb{Z}[F])$
is the obstruction class: if $\delta s]\neq 0$ then there is no global section compatible with the given local data.
In the Abramsky–Brandenburger sheaf-theoretic framework, an empirical model is a compatible family of probability distributions on maximal contexts, and contextuality is the non-existence of a global section explaining these marginals [18]. Abramsky, Mansfield and Barbosa define a Čech-cohomological obstruction class (built from an abelian presheaf derived from the support of the model) which vanishes whenever a global section exists; hence non-vanishing provides a robust sufficient witness of contextuality [34, 35].
Interpretation in Physics
(a) Classical Physics
If all local data glue globally:
$\check{H}^{1}(\mathcal{C},F)=0.$
Examples:
  • classical EM fields ($\mathbf{E},\mathbf{B}$),
  • classical trajectories,
  • classical probability distributions.
(b) Quantum or Contextual Physics
If gluing fails:
$\check{H}^{1}(\mathcal{C},F)\neq 0.$
This signals:
  • contextuality,
  • nonclassical phase structure,
  • interference phenomena,
  • nonexistence of global truth-values.
An example is the Kochen-Specker obstruction [36, 37].
12 Appendix B: A Seven-valued Contextual Logic
Ghose and Patra [24] have proposed a many-valued and contextual logic. In the GP scheme, the seven values are understood as describing how the truth of a proposition $P$ varies across incompatible measurement contexts. For example, the value corresponding to “true-and-false” does not assert that $P$ is both true and false in the same context; rather, it asserts that $P$ is true in some context $C_{1}$ and false in another, incompatible context $C_{2}$. The combined values record patterns of variation across contexts.
Using the quantifier $\forall$, GP have shown that the basic modes can be formally written as (i) $\forall x\,[\phi(x)\rightarrow p(x)]$; (ii) $\forall x\,[\phi(x)\rightarrow\neg p(x)]$; (iii) $\forall x\,[\phi(x)\rightarrow q(x)]$, $x$ standing for a variable (a placeholder) which ranges over the domain of a system (like pots), $\phi$ for a well formed formula that specifies some condition (like for example ‘baked’), $p$ for some predicate (like say ‘red’) and $q$ for the predicate avaktavyam. An ‘example’ of the first of these three in plain English would be: for all $x$ (say clay pots) the condition $\phi(x)$ (say ‘baked’) implies that the pot is red.
The other four compounds may be written as
(iv) $\forall x\,[\phi(x)\rightarrow p(x)\land\phi^{\prime}(x)\rightarrow\neg p(x)]\land\neg[\phi(x)\leftrightarrow\phi^{\prime}(x)]$,
(v) $\forall x\,[\phi(x)\rightarrow p(x)\land\phi^{\prime}(x)\rightarrow q(x)]\land\neg[\phi(x)\leftrightarrow\phi^{\prime}(x)]$,
(vi) $\forall x\,[\phi(x)\rightarrow\neg p(x)\land\phi^{\prime}(x)\rightarrow q(x)]\land\neg[\phi(x)\leftrightarrow\phi^{\prime}(x)]$,
(vii) $\forall x\,[\phi(x)\rightarrow p(x)\land\phi^{\prime}(x)\rightarrow\neg p(x)\land\phi^{\prime\prime}(x)\rightarrow q(x)]\land\neg[\phi(x)\leftrightarrow\phi^{\prime}(x)]\land\neg[\phi^{\prime}(x)\leftrightarrow\phi^{\prime\prime}(x)]\land\neg[\phi(x)\leftrightarrow\phi^{\prime\prime}(x)]$.
Written in this formal way, the seven predications are self-consistent as they hold under mutually exclusive conditions.
Acknowledgements
The author acknowledges the use of ChatGPT (OpenAI) for language polishing and assistance with formatting. The author takes full responsibility for the scientific content, interpretations, and any remaining errors.